# Efficient detection of entanglement by stimulated disentanglement

### v1

@@ 4.2 Experimental Methodology
After compensating the dispersion, we also need to stabilize the global phase of the SPDC spectrum relative to the pump. We use for that purpose a Pound-Drever-Hall (PDH) feedback loop, where the phase modulation of the PDH loop is later also used for the lock-in detection of the SFG signal. Conceptually, one could utilize the Intensity of the entangled bi-photons at the output of the SU(1,1) interference to retrieve the phase between the photon-pairs and the pump to lock the interferometer to constructive (PDC) or destructive (SFG) interference. However, the SPDC intensity at the single photon regime is too weak and its SNR is low, leading to an unstable and fragile phase lock. Thus, to enhance the measurement SNR and improve the stability and robustness of the phase lock, we seed the SU(1,1) input with a specific signal frequency (1560 nm), stimulating the PDC to generate a strong matching idler at $\lambda_{i}=2035$ nm, whose intensity is sufficient to serve as a reliable reference for phase locking with a satisfactory SNR. The phase lock is implemented using a fast piezoelectric actuator on one of the mirrors in the spectral shaper [45].
@@ 5.1 The non-classical response to loss
To gain deeper insight into the detection dynamics of our coherent SFG method and the relations between the interfering SFG contributions - spontaneous and stimulated, we investigated the response of the SFG signal to SPDC loss, regarding the theoretical prediction of Eq. 12. SPDC loss affects both photons of a pair, leading to a unique quantum situation, where although the flux of individual photons is reduced by $T$ the bi-photon rate is reduced quadratically, as the probability of both photons surviving the SPDC loss is $T^{2}$.
@@
Funding This work was supported by the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101178170 and from the Israel Science Foundation under grant No. 2208/24.
@@
Acknowledgments The section title should not follow the numbering scheme of the body of the paper. Additional information crediting individuals who contributed to the work being reported, clarifying who received funding from a particular source, or other information that does not fit the criteria for the funding block may also be included; for example, “K. Flockhart thanks the National Science Foundation for help identifying collaborators for this work.”

### v2

@@ 4.2 Experimental Methodology
After compensating the dispersion, we also need to stabilize the global phase of the SPDC spectrum relative to the pump. We use for that purpose a Pound-Drever-Hall (PDH) feedback loop, where the phase modulation of the PDH loop is later also used for the lock-in detection of the SFG signal. Conceptually, one could utilize the Intensity of the entangled biphotons at the output of the SU(1,1) interference to retrieve the phase between the photon-pairs and the pump to lock the interferometer to constructive (PDC) or destructive (SFG) interference. However, the SPDC intensity at the single photon regime is too weak and its SNR is low, leading to an unstable and fragile phase lock. Thus, to enhance the measurement SNR and improve the stability and robustness of the phase lock, we seed the SU(1,1) input with a specific signal frequency (1560 nm), stimulating the PDC to generate a strong matching idler at $\lambda_{i}=2035$ nm, whose intensity is sufficient to serve as a reliable reference for phase locking with a satisfactory SNR. The phase lock is implemented using a fast piezoelectric actuator on one of the mirrors in the spectral shaper [45].
@@ 5.1 The non-classical response to loss
To gain deeper insight into the detection dynamics of our coherent SFG method and the relations between the interfering SFG contributions - spontaneous and stimulated, we investigated the response of the SFG signal to SPDC loss, regarding the theoretical prediction of Eq. 12. SPDC loss affects both photons of a pair, leading to a unique quantum situation, where although the flux of individual photons is reduced by $T$ the biphoton rate is reduced quadratically, as the probability of both photons surviving the SPDC loss is $T^{2}$.
@@
Funding This project was funded by the Israel Innovation Authority and Elta Systems Ltd. under project No. 70002 and under a Danish-Israeli Eureka project. Funding is acknowledged also from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101178170 and from the Israel Science Foundation under grant No. 2208/24.
@@
Acknowledgments We wish to thank Ruben Faibish, David Karasik and Nissan Maskil for helpful discussions.